OpenStudy (anonymous):

Solve: -(3x - 2) = -4(x + 1) - 4 a. -2 b. 2 c. 10 d. -10

OpenStudy (anonymous):

@whpalmer4 ?

OpenStudy (whpalmer4):

\[-(3x-2) = -4(x+1)-4\]Use the distributive property to remove the parentheses. Hint: \[-(x-1) = -1(x-1)\]

OpenStudy (anonymous):

D

OpenStudy (whpalmer4):

Distributive property, in case you don't remember it: \[a(b+c) = a*b + a*c\]

OpenStudy (anonymous):

is d correct though?

OpenStudy (whpalmer4):

You'll learn best if you work it out yourself. Go through the steps and I'll check your work.

OpenStudy (whpalmer4):

\[-(3x-2) = \]